OpenStudy (anonymous):

graph the line y= - 2/5 x + 1 using the slope and the y- intercept

OpenStudy (anonymous):

Hey!

OpenStudy (anonymous):

help

OpenStudy (lukebluefive):

The y-intercept will be 1, since setting the x in the equation equal to 0 gives: y = 0 + 1 Then, use the slope to make another point two units over and five units down, due to the equation of a slope: \[slope = \frac{rise}{run}\]

OpenStudy (anonymous):

The general procedure is to start at the y intercept on your graph, draw a point there, and then traverse rise and run of the slope to find another point then draw a line through them. For example for \[y=\frac{2}{3}x+5\] the y intercept is 5. Therefore there is a point at (0,5). Draw a point there. Then FROM THAT POINT move up 2 (the numerator of the slope) and to the right 3 and draw another point. Then draw a line through the two points put arrows on the line, make sure you've labeled everything and you are don.

OpenStudy (jusaquikie):

y=mx+b b is where the line crosses the y axis, and m is the slope, which means if you have an x of 1 you will have a y of -2/5(1) +1